Thelma is planning for her sons college education to begin five years from today. She estimates the yearly tuition, books, and living expenses to be $5,000 per year for a four-year degree, assuming the expenses incur only at the end of the year. How much must Thelma deposit today, at an interest rate of 8 percent, for her son to be able to withdraw $5,000 per year for four years of rollege? 16 $11,270 $13,620 $20,000 $39,520
Answer #1

Let us assume that amount invested be x

rate of interest = 8% = 0.08

Hence, amount accumulated after 5 years = x*1.085

Now, the son withdraws 5000 per year for 4 years

The Present value of these withdrawals 5 years from now = 5000/1.08 + 5000/1.082 + 5000/1.083 + 5000/1.084= 16560.63

Now, amount accumulated after 5 years = value of 4 withdrawals at Year 5

=> 1.085x = 16560.63

=> x = $11270

Hence, option (a) $11270 is the correct option